OpenStudy (anonymous):

Solve. 3x2 + 12x = 3

OpenStudy (anonymous):

Do you intend the following?\[3x2+12x=3\]...or in other words...\[6x+12x=3\]

OpenStudy (mertsj):

Divide by 3 first.

OpenStudy (mertsj):

x^2+4x-1=0

OpenStudy (anonymous):

Or \[3x^2+12x=3\]?

OpenStudy (anonymous):

you can use quadratic equation, or factor. 3x2 + 12x -3 = 0 ==> x2 + 4x -1 = 0 since there is no easy factoring method, u should use the quadratic equation.

OpenStudy (anonymous):

where it is in the form ax2 + bx + c = 0 and the 2 x values = \[(-b \pm \sqrt{b^{2}- 4ac})/2a\]

OpenStudy (mertsj):